Circular No. 7627

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN AQUILA G. M. Hurst, The Astronomer, Basingstoke, England, communicates:

“M. Collins, Everton, Bedfordshire, has reported his detection of a variable object in Aquila during photography with a 135-mm telephoto lens on Kodak 2415 film for the U.K. Nova/Supernova Patrol. The object was recorded on May 11.988 UT at mag 10.9. An image by Collins from 2000 Nov. 26.8 failed to show the object (limiting mag 12.8), which was also absent from a master patrol image obtained on 1997 Aug. 25.9 (limiting mag 13). A patrol image from 2001 Apr. 25.1 may show the object near mag 12.1, but as the Milky Way field is crowded this is rather uncertain. In response to an appeal, N. D. James, Chelmsford (0.30-m reflector), obtained a 30-s CCD frame of the field on May 16.966 UT and found the object at V = 12.6 (GSC system) in the measured position α = 19^{h}07^{m}28^{s}.41, δ = +11^{o}44′45′′.8 (equinox 2000.0; 15 GSC stars, mean error 0′′.24 in α, 0′′.15 in δ). The nearest star (R = 16.2) in the USNO A2.0 catalogue has position end figures 28^{s}.34, 52′′.1.”

COMET C/2001 A2 (LINEAR) E. Jehin, A. Jaunsen, H. Boehnhardt, M. Kiekebusch, H. Nunez, R.

Amestica, C. Herrera, J. Navarete, F. Delgado and R. M. West, European Southern Observatory, report: “Images of comet C/2001 A2 have been obtained using the 8.2-m Very Large Telescopes Melipal and Yepun with the Nasmyth and Cassegrain test cameras, respectively. On May 14.98 UT two components were seen in R-band images, the eastern, tailward one (component A) ∼ 1 mag fainter than component B (within an aperture of 1′′.3) at a separation of 12′′.6 in p.a. 105^{o}. Both components had individual comae elongated approximately in the antisolar direction. Component B showed sunward-extended isophotes in the very inner part of the coma. On May 16.98 UT the distance between the components had increased to 14′′.6 (in the same position angle). In addition, the sunward fragment appeared to have split into two components with a separation of 1′′ in p.a. 135^{o}–315^{o}; these components were of about the same brightness (in R) and surrounded by a joint coma. V -band images revealed very extended isophotes perpendicular to the separation direction of this new pair. This could indicate the presence of a large amount of gas in the coma in addition to the dust.”

J. Broughton, Reedy Creek, Queensland (0.25-m Schmidt-Cassegrain), reports further CCD astrometry on May 14.4 and 16.3 UT, noting that on the latter occasion component A was at least two magnitudes fainter than and was separated from component B by about 14′′ in p.a. 102^{o}.

2001 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden